# LLMs for Knowledge Graph Construction and Reasoning: Recent Capabilities and Future Opportunities

### v2

@@ Abstract
This paper presents an exhaustive quantitative and qualitative evaluation of Large Language Models (LLMs) for Knowledge Graph (KG) construction and reasoning. We engage in experiments across eight diverse datasets, focusing on four representative tasks encompassing entity and relation extraction, event extraction, link prediction, and question-answering, thereby thoroughly exploring LLMs’ performance in the domain of construction and inference. Empirically, our findings suggest that LLMs, represented by GPT-4, are more suited as inference assistants rather than few-shot information extractors. Specifically, while GPT-4 exhibits good performance in tasks related to KG construction, it excels further in reasoning tasks, surpassing fine-tuned models in certain cases. Moreover, our investigation extends to the potential generalization ability of LLMs for information extraction, leading to the proposition of a Virtual Knowledge Extraction task and the development of the corresponding VINE dataset. Based on these empirical findings, we further propose AutoKG, a multi-agent-based approach employing LLMs and external sources for KG construction and reasoning. We anticipate that this research can provide invaluable insights for future undertakings in the field of knowledge graphs  [^1] The code and datasets are in https://github.com/zjunlp/AutoKG..
LLMs for Knowledge Graph Construction and Reasoning:
Recent Capabilities and Future Opportunities
[…]
@@ 1 Introduction
Knowledge Graph (KG) is a semantic network comprising entities, concepts, and relations (Cai et al., 2022; Zhu et al., 2022; Liang et al., 2022; Chen et al., 2023; Pan et al., 2023b, a), which can catalyse applications across various scenarios. Constructing KGs (Ye et al., 2022b) typically involves multiple tasks such as Named Entity Recognition (NER) (Chiu and Nichols, 2016), Relation Extraction (RE) (Zeng et al., 2015; Chen et al., 2022), Event Extraction (EE) (Chen et al., 2015; Deng et al., 2020), and Entity Linking (EL) (Shen et al., 2015). Additionally, Link Prediction (LP) (Zhang et al., 2018; Rossi et al., 2021) is a crucial step for KG reasoning, essential for understanding constructed KGs. These KGs also hold a central position in Question Answering (QA) tasks (Karpukhin et al., 2020; Zhu et al., 2021), especially in conducting inference based on question context, involving the construction and application of relation subgraphs. This paper focuses on empirically investigating the potential applicability of LLMs (Liu et al., 2023; Shakarian et al., 2023; Lai et al., 2023; Zhao et al., 2023b), exemplified by ChatGPT and GPT-4 (OpenAI, 2023). By comprehending the fundamental capabilities of LLMs, our study further delves into potential future directions.
@@
Entity and Relation Extraction and Event Extraction serve as foundational elements for constructing knowledge graphs, facilitating the refinement of a wealth of entity, relation, and event information. Meanwhile, Link Prediction, as a core task of KG reasoning, aims to uncover latent relationships between entities, thereby enriching the knowledge graph. Additionally, we further explore the application of LLMs in knowledge-based Question Answering tasks to gain a comprehensive understanding of their inferential skills. Given these considerations, we select these tasks as representatives for evaluating both the construction and reasoning of KGs. As illustrated in Figure 1, our initial investigation targets the zero-shot and one-shot abilities of large language models across the aforementioned tasks. This analysis serves to assess the potential usage of such models in the field of knowledge graphs. The empirical findings reveal that LLMs like GPT-4 exhibit limited effectiveness as a few-shot information extractor, yet demonstrate considerable proficiency as an inference assistant.
Generalizability Analysis.
@@
To delve deeper into the behavior of LLMs in information extraction tasks, we devise a unique task termed Virtual Knowledge Extraction. This undertaking aims to discern whether the observed performance enhancements on these tasks are attributed to the extensive internal knowledge repositories of LLMs or to their potent generalization capabilities facilitated by instruction tuning and Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017). And our experiments on a newly constructed dataset, VINE, indicate that large language models like GPT-4 can acquire new knowledge from instructions and effectively execute extraction tasks, thereby affording a more nuanced understanding of large models to a certain extent.
Future Opportunities.
@@
In light of the preceding experiments, we further examine prospective directions for knowledge graphs. Given the remarkable generalization capabilities of large models, we opt to employ them to aid in the construction of KG. Compared to smaller models, these LLMs mitigate potential resource wastage and demonstrate notable adaptability in novel or data-scarce situations. However, it’s important to recognize their strong dependence on prompt engineering and the inherent limitations of their knowledge cutoff. Consequently, researchers are exploring interactive mechanisms that allow LLMs to access and leverage external resources, aiming to enhance their performance further (Wang et al., 2023b).
@@
On this basis, we introduce the concept of AutoKG - autonomous KG construction and reasoning via multi-agent communication. In this framework, the human role is diminished, with multiple communicative agents each playing their respective roles. These agents interact with external sources, collaboratively accomplishing the task. We summarize our contributions as follows:
@@ 2 Recent Capabilities of LLMs for KG Construction and Reasoning
The release of large language models like GPT-4, recognized for their remarkable general capabilities, has been considered by researchers as the spark of artificial general intelligence (AGI) (Bubeck et al., 2023). To facilitate an in-depth understanding of their performance in KG-related tasks, a series of evaluations are conducted. §2.1 introduces the evaluation principles, followed by a detailed analysis in §2.2 on the performance of LLMs in the construction and reasoning tasks, highlighting variations across different datasets and domains. Moreover, §2.3 delves into the reasons underlying the subpar performance of LLMs in certain tasks. And finally, §2.4 discusses whether the models’ performance is genuinely indicative of generalization abilities or influenced by inherent advantages of the knowledge base.
@@ 2.1 Evaluation Principle
In this study, we conduct a comprehensive assessment of LLMs, represented by GPT-4, and specifically analyze the performance disparities and enhancements between GPT-4 and other models in the GPT series, such as ChatGPT. A primary area of investigation is the models’ performance in zero-shot and one-shot tasks, as these tasks illuminate the models’ generalization capabilities under data-limited conditions. Utilizing the evaluation results, our objective is to explore the reasons behind the models’ exemplary performance in specific tasks and identify potential areas of improvement. Ultimately, our goal is to derive valuable insights for future advancements in such models.
@@ 2.2.1 Settings
Datasets.
During the task of Entity and Relation Extraction, Event Extraction, we employ DuIE2.0 (Li et al., 2019), SciERC (Luan et al., 2018), Re-TACRED (Stoica et al., 2021), and MAVEN (Wang et al., 2020) datasets. For Link Prediction, we utilized FB15K-237 (Toutanova et al., 2015) and ATOMIC 2020 (Hwang et al., 2021a) datasets. Finally, FreebaseQA (Jiang et al., 2019) and MetaQA (Zhang et al., 2018) datasets are used in the Question Answering task. The dataset used is described in detail in Appendix B.
@@ 2.2.2 Overall Results
Entity and Relation Extraction.
We conduct experiments on DuIE2.0, Re-TACRED, and SciERC, each involving 20 samples in the test/valid sets, encompassing all types of relationships present within the datasets. Here we use PaddleNLP LIC2021 IE   [^2] https://github.com/PaddlePaddle/PaddleNLP/tree/develop/examples/information_extraction/DuIE., PL-Marker (Ye et al., 2022a) and EXOBRAIN (Park and Kim, 2021) as baselines on each dataset, respectively. Concurrently, for evaluation purposes, the results are reported utilizing the standard micro F1 score. As shown in Table 1, GPT-4 performs relatively well in both zero-shot and one-shot manners compared to ChatGPT, even though its performance has not yet surpassed that of fully supervised small models.
@@
Figure 3: An example of MAVEN by GPT-4.
Model Knowledge Graph Construction
DuIE2.0 Re-TACRED SciERC MAVEN
Fine-Tuned SOTA 69.42 91.4 53.2 68.8
@@
text-davinci-003 11.43 9.8 4.0 30.0
ChatGPT 10.26 15.2 4.4 26.5
GPT-4 31.03 15.5 7.2 34.2
@@
text-davinci-003 30.63 12.8 4.8 25.0
ChatGPT 25.86 14.2 5.3 34.1
GPT-4 41.91 22.5 9.1 30.4
Table 1: KG Construction tasks (F1 score).
Event Extraction.
For simplification, we conduct event detection experiments on 20 random samples from MAVEN, encompassing all event types. Using the F-score metric, GPT-4’s performance is benchmarked against the existing state-of-the-art (SOTA) (Wang et al., 2022a) model, as well as to other models in the GPT family. Based on our results, GPT-4 shows inconsistent superiority over the SOTA, with both GPT-4 and ChatGPT outperforming each other in different scenarios.
$\bullet$ Zero-shot As shown in Table 1, GPT-4 outperforms ChatGPT. For the sentence “Now an established member of the line-up, he agreed to sing it more often.”, ChatGPT generates the result Becoming_normal-_\__a_normal-_\__member, while GPT-4 identifies two more: Agree_normal-_\__or_normal-_\__refuse_normal-_\__to_normal-_\__act, Performing. It is worth noting that in this experiment, ChatGPT frequently provides answers with a single event type. In contrast, GPT-4’s ability to grasp complex contextual information enables it to identify multiple event types within these sentences.
$\bullet$ One-shot In this configuration, ChatGPT’s performance improves notably, while GPT-4 experiences a slight decline. Figure 3 illustrates that GPT-4 incorrectly identifies five event types where the correct answers are Process_normal-_\__end and Come_normal-_\__together. Despite detecting underlying ranking and comparison information, GPT-4 misses the trigger words final and host. Simultaneously, we observe that under one-shot setup, GPT-4 tends to produce a higher number of erroneous responses when it is unable to identify the correct ones. We theorize this could stem from implicit type indications of the dataset.
Link Prediction.
Task link prediction involves experiments on two distinct datasets, FB15k-237 and ATOMIC2020. The former is a random sample set comprising 25 instances, whereas the latter encompasses 23 instances on behalf of all possible relations. Among various approaches, the best performing fine-tuned models are C-LMKE (BERT-base) (Wang et al., 2022b) and COMET (BART) (Hwang et al., 2021b) for each.
@@
Figure 4: Examples of task LP and QA.
Model Knowledge Graph Reasoning
FB15K-237 ATOMIC2020 FreebaseQA MetaQA
Fine-Tuned SOTA 32.4 46.9 79.0 100
Zero-shot
text-davinci-003 16.0 15.1 95.0 33.9
ChatGPT 24.0 10.6 95.0 52.7
GPT-4 32.0 16.3 95.0 63.8
One-shot
text-davinci-003 32.0 14.1 95.0 49.5
ChatGPT 32.0 11.1 95.0 50.0
GPT-4 40.0 19.1 95.0 56.0
Table 2: Link Prediction (hits@1 / bleu1) and Question Answering (AnswerExactMatch).
Question Answering.
We conduct the evaluation using two prevalent Knowledge Base Question Answering datasets, FreebaseQA and MetaQA, with 20 random instances sampled from each. In MetaQA, we sample proportional to their dataset representation. Yu et al. (2022) and  Madani and Joseph (2023) represent the SOTA models employed. And for both datasets, AnswerExactMatch is adopted as the metric of evaluation.
@@
Our experiments on KG construction and reasoning reveal that LLMs exhibit superior reasoning skills compared to their construction capabilities. Given the absence of more refined evaluation standards, we assess the comparative capabilities of LLMs in these tasks by measuring the performance differential between LLMs and the current SOTA methodologies. Despite the exemplary performance of LLMs, they do not surpass the current state-of-the-art models in KG construction under zero-shot and one-shot settings, indicating limitations in extracting information from sparse data. Conversely, all LLMs in one-shot, and GPT-4 in zero-shot, match or near SOTA performance on the FreebaseQA and FB15K-237 datasets. Moreover, they exhibit relatively good performance across the remaining datasets, which underscores their adaptability in KG reasoning tasks as well. The intrinsic complexity of KG construction tasks may account for this discrepancy in performance. Furthermore, the robust reasoning performance of LLMs might be attributed to their exposure to relevant knowledge during pre-training.
@@
In our study, we evaluate the performance of large language models, exemplified by GPT-4, across diverse knowledge domains, ensuring a balanced assessment in both generic and specialized contexts. The chosen benchmarks, SciERC and Re-TACRED, represent scientific and general domains, respectively. While Re-TACRED exhibits a broader range of relation types compared to the seven in SciERC, both GPT-4 and ChatGPT underperform on the specialized SciERC dataset, indicating their limitations in domain-specific data. Interestingly, GPT-4’s performance boost on SciERC is less pronounced than on Re-TACRED when given one demonstration. We hypothesize that the subpar performance on specialized datasets may stem from these models being predominantly trained on vast general corpora, thereby lacking sufficient domain-specific expertise.

### v3

@@ Abstract
This paper presents an exhaustive quantitative and qualitative evaluation of Large Language Models (LLMs) for Knowledge Graph (KG) construction and reasoning. We engage in experiments across eight diverse datasets, focusing on four representative tasks encompassing entity and relation extraction, event extraction, link prediction, and question-answering, thereby thoroughly exploring LLMs’ performance in the domain of construction and inference. Empirically, our findings suggest that LLMs, represented by GPT-4, are more suited as inference assistants rather than few-shot information extractors. Specifically, while GPT-4 exhibits good performance in tasks related to KG construction, it excels further in reasoning tasks, surpassing fine-tuned models in certain cases. Moreover, our investigation extends to the potential generalization ability of LLMs for information extraction, leading to the proposition of a Virtual Knowledge Extraction task and the development of the corresponding VINE dataset. Based on these empirical findings, we further propose AutoKG, a multi-agent-based approach employing LLMs and external sources for KG construction and reasoning. We anticipate that this research can provide invaluable insights for future undertakings in the field of knowledge graphs.
keywords:
Knowledge Graph, Information Extraction, GPT-4, Large Language Model
@@ 1 Introduction
Knowledge Graph (KG) is a semantic network comprising entities, concepts, and relations [1, 2, 3, 4, 5, 6], which can catalyse applications across various scenarios. Constructing KGs [7, 8] typically involves multiple tasks such as Named Entity Recognition (NER) [9, 10], Relation Extraction (RE) [11, 12], Event Extraction (EE) [13, 14], and Entity Linking (EL) [15]. Additionally, Link Prediction (LP) [16, 17] is a crucial step for KG reasoning, essential for understanding constructed KGs. These KGs also hold a central position in Question Answering (QA) tasks [18, 19], especially in conducting inference based on question context, involving the construction and application of relation subgraphs. This paper empirically investigates the potential applicability of LLMs in the KG domain, taking ChatGPT and GPT-4 [20] as examples. The research begins with an examination of the fundamental capabilities of LLMs [21, 22, 23, 24], progressing to explore possible future developments, aiming to enhance our understanding of LLMs and introduce new perspectives and methods to the field of knowledge graphs.
@@
Recent Capabilities. Entity and Relation Extraction, along with Event Extraction, are pivotal for Knowledge Graph (KG) construction tasks [25, 26, 27, 28]. They play a critical role in organizing vast amounts of entity, relation, and event data into structured representations, forming the foundational elements that underpin the construction and enrichment of KGs. Meanwhile, Link Prediction, as a core task of KG reasoning [29], aims to uncover latent relationships between entities, thereby enriching the knowledge graph. Additionally, we delve into the utilization of LLMs in knowledge-based Question Answering tasks [30, 31] to gain a thorough insight into their reasoning capabilities. Given these considerations, we select these tasks as representatives for evaluating both the construction and reasoning of KGs. As illustrated in Figure 1, our initial investigation targets the zero-shot and one-shot abilities of large language models across the aforementioned tasks. This analysis serves to assess the potential usage of such models in the field of knowledge graphs. The empirical findings reveal that LLMs like GPT-4 exhibit limited effectiveness as a few-shot information extractor, yet demonstrate considerable proficiency as an inference assistant.
@@
Generalizability Analysis. To delve deeper into the behavior of LLMs in information extraction tasks, we devise a unique task termed Virtual Knowledge Extraction, targeting LLMs’ ability to generalize and extract unfamiliar knowledge. This undertaking aims to discern whether the observed performance enhancements on these tasks are attributed to the extensive internal knowledge repositories of LLMs or to their potent generalization capabilities facilitated by instruction tuning [32] and Reinforcement Learning from Human Feedback (RLHF) [33]. And our experiments on a newly constructed dataset, VINE, indicate that large language models like GPT-4 can acquire new knowledge from instructions and effectively execute extraction tasks, thereby affording a more nuanced understanding of large models to a certain extent.
@@
Future Opportunities. In light of the preceding experiments, we further examine prospective directions for knowledge graphs. Given the remarkable generalization capabilities of large models [34, 35], we opt to employ them to aid in the construction of KG. Compared to smaller models, these LLMs mitigate potential resource wastage and demonstrate notable adaptability in novel or data-scarce situations. However, it’s important to recognize their strong dependence on prompt engineering [36] and the inherent limitations of their knowledge cutoff. Consequently, researchers are exploring interactive mechanisms that allow LLMs to access and leverage external resources, aiming to enhance their performance further [37].
@@
On this basis, we introduce the concept of AutoKG - autonomous KG construction and reasoning via multi-agent communication. In this framework, the human role is diminished, with multiple communicative agents each playing their respective roles. These agents interact with external sources, collaboratively accomplishing the task. We summarize our contributions as follows  [^1] The code and datasets are in https://github.com/zjunlp/AutoKG.:
@@ 2 Recent Capabilities of LLMs for KG Construction and Reasoning
The release of large language models like GPT-4, recognized for their remarkable general capabilities, has been considered by researchers as the spark of artificial general intelligence (AGI) [38]. To facilitate an in-depth understanding of their performance in KG-related tasks, a series of evaluations are conducted. §2.1 introduces the evaluation principles, followed by a detailed analysis in §2.2 on the performance of LLMs in the construction and reasoning tasks, highlighting variations across different datasets and domains. Moreover, §2.3 delves into the reasons underlying the subpar performance of LLMs in certain tasks. And finally, §2.4 discusses whether the models’ performance is genuinely indicative of generalization abilities or influenced by inherent advantages of the knowledge base.
@@ 2.1 Evaluation Principle
In this study, we conduct a comprehensive assessment of LLMs, represented by GPT-4, and specifically analyze the performance disparities and enhancements between GPT-4 and other models in the GPT series, such as ChatGPT. A primary area of investigation is the models’ performance in zero-shot and one-shot tasks, as these tasks illuminate the models’ generalization capabilities under data-limited conditions. Given that some experiments in our study rely on randomly sampled subsets of datasets, it is important to note that there may be inherent variability in the results due to this sampling approach. We deliberately choose zero-shot and one-shot tasks over those requiring more examples, as they better test the models’ adaptability and practical application in scenarios with sparse data. The experimental prompt used is detailed in Appendix D. Utilizing the evaluation results, our objective is to explore the reasons behind the models’ exemplary performance in specific tasks and identify potential areas of improvement. Ultimately, our goal is to derive valuable insights for future advancements in such models.
@@ 2.2.1 Settings
Datasets. During the task of Entity and Relation Extraction, Event Extraction, we employ DuIE2.0 [39], SciERC [40], Re-TACRED [41], and MAVEN [42] datasets. For Link Prediction, we utilized FB15K-237 [43] and ATOMIC 2020 [44] datasets. Finally, FreebaseQA [45] and MetaQA [16] datasets are used in the Question Answering task. The dataset used is described in detail in Appendix B.
@@ 2.2.2 Overall Results
Entity and Relation Extraction. We conduct experiments on DuIE2.0, Re-TACRED, and SciERC, each involving 20 samples in the test/valid sets, encompassing all types of relationships present within the datasets. Here we use PaddleNLP LIC2021 IE   [^2] https://github.com/PaddlePaddle/PaddleNLP/tree/develop/examples/information_extraction/DuIE., PL-Marker [46] and EXOBRAIN [47] as baselines on each dataset, respectively. Concurrently, for evaluation purposes, the results are reported utilizing the standard micro F1 score. As shown in Table 1, GPT-4 performs relatively well in both zero-shot and one-shot manners compared to ChatGPT, even though its performance has not yet surpassed that of fully supervised small models.
@@
Figure 3: Here are examples of task Event Extraction, Link Prediction and Question Answering.
Event Extraction. For simplification, we conduct event detection experiments on 20 random samples from MAVEN, encompassing all event types. Using the F-score metric, GPT-4’s performance is benchmarked against the existing state-of-the-art (SOTA) [48] model, as well as to other models in the GPT family. Based on our results, GPT-4 shows inconsistent superiority over the SOTA, with both GPT-4 and ChatGPT outperforming each other in different scenarios.
$\bullet$ Zero-shot As shown in Table 1, GPT-4 outperforms ChatGPT. For the sentence “Now an established member of the line-up, he agreed to sing it more often.”, ChatGPT generates the result Becoming__\__a__\__member, while GPT-4 identifies two more: Agree__\__or__\__refuse__\__to__\__act, Performing. It is worth noting that in this experiment, ChatGPT frequently provides answers with a single event type. In contrast, GPT-4’s ability to grasp complex contextual information enables it to identify multiple event types within these sentences.
$\bullet$ One-shot In this configuration, ChatGPT’s performance improves notably, while GPT-4 experiences a slight decline. Figure 3 illustrates that GPT-4 incorrectly identifies five event types where the correct answers are Process__\__end and Come__\__together. Despite detecting underlying ranking and comparison information, GPT-4 misses the trigger words final and host. Simultaneously, we observe that under one-shot setup, GPT-4 tends to produce a higher number of erroneous responses when it is unable to identify the correct ones. We theorize this could stem from implicit type indications of the dataset.
Table 1: KG Construction and KG Reasoning tasks.
Model Knowledge Graph Construction Knowledge Graph Reasoning
DuIE2.0 Re-TACRED SciERC MAVEN FB15K-237 ATOMIC2020 FreebaseQA MetaQA
Fine-Tuned SOTA 69.42 91.4 53.2 68.8 32.4 46.9 79.0 100
@@
text-davinci-003 11.43 9.8 4.0 30.0 16.0 15.1 95.0 33.9
ChatGPT 10.26 15.2 4.4 26.5 24.0 10.6 95.0 52.7
GPT-4 31.03 15.5 7.2 34.2 32.0 16.3 95.0 63.8
@@
text-davinci-003 30.63 12.8 4.8 25.0 32.0 14.1 95.0 49.5
ChatGPT 25.86 14.2 5.3 34.1 32.0 11.1 95.0 50.0
GPT-4 41.91 22.5 9.1 30.4 40.0 19.1 95.0 56.0
Link Prediction. Task link prediction involves experiments on two distinct datasets, FB15k-237 and ATOMIC2020. The former is a random sample set comprising 25 instances, whereas the latter encompasses 23 instances on behalf of all possible relations. Among various approaches, the best performing fine-tuned models are C-LMKE (BERT-base) [49] and COMET (BART) [50] for each.
@@
Question Answering. We conduct the evaluation using two prevalent Knowledge Base Question Answering datasets, FreebaseQA and MetaQA, with 20 random instances sampled from each. In MetaQA, we sample proportional to their dataset representation. Yu et al. [51] and  Madani and Joseph [52] represent the SOTA models employed. And for both datasets, AnswerExactMatch is adopted as the metric of evaluation.
@@
Our experiments on KG construction and reasoning reveal that LLMs exhibit superior reasoning skills compared to their construction capabilities. Given the challenge of quantifying reasoning and construction abilities, we assess the comparative capabilities of LLMs in these tasks by measuring the performance differential between LLMs and the current SOTA methodologies. Larger performance disparities indicate poorer performance. Despite the exemplary performance of LLMs, they do not surpass the current state-of-the-art models in KG construction under zero-shot and one-shot settings, indicating limitations in extracting information from sparse data. Conversely, all LLMs in one-shot, and GPT-4 in zero-shot, match or near SOTA performance on the FreebaseQA and FB15K-237 datasets. Moreover, they exhibit relatively good performance across the remaining datasets, which underscores their adaptability in KG reasoning tasks as well. The intrinsic complexity of KG construction tasks may account for this discrepancy in performance. Furthermore, the robust reasoning performance of LLMs might be attributed to their exposure to relevant knowledge during pre-training.
@@
In our study, we evaluate the performance of large language models, exemplified by GPT-4, across diverse knowledge domains, ensuring a balanced assessment in both generic and specialized contexts. We employed a consistent method of evaluating relative task capabilities, similar to the performance disparity assessment described in §2.2.3. The chosen benchmarks, SciERC and Re-TACRED, represent scientific and general domains, respectively. While Re-TACRED exhibits a broader range of relation types compared to the seven in SciERC, both GPT-4 and ChatGPT underperform on the specialized SciERC dataset, indicating their limitations in domain-specific data. Interestingly, GPT-4’s performance boost on SciERC is less pronounced than on Re-TACRED when given one demonstration. Specifically, during our experiments, we note challenges in LLMs’ recognition and understanding of specialized terms within the SciERC dataset. We hypothesize that the subpar performance on specialized datasets may stem from these models being predominantly trained on vast general corpora, thereby lacking sufficient domain-specific expertise.